# WebRL: Training LLM Web Agents via Self-Evolving Online Curriculum Reinforcement Learning

### v1

@@ Abstract
Figure 1: (a) Compared with all proprietary and open-sourced LLMs, GLM-4-9B with WebRL achieves the best results. (b) The performance of GLM-4-9B on WebArena-Lite (Zhou et al., 2023a; Liu et al., 2024), trained using WebRL, shows significant improvement over other baselines across all five evaluated websites.
@@ 1 Introduction
Large language models (LLMs) have exhibited not only superior comprehension of human language, commonsense reasoning, and knowledge acquisition, but also significant potential in complex planning and logical reasoning, indicating their promising trajectory towards serving as autonomous LLM agents (Wang et al., 2023; Liu et al., 2023a). A diverse array of applications for LLM agents has proliferated, encompassing domains such as code generation (Jimenez et al., 2024), database manipulation (Zhou et al., 2023b; Gu et al., 2024), and graphical user interface (GUI) interaction (Rawles et al., 2024; Yang et al., 2023; Xie et al., 2024). Among these, web agents powered by LLMs (Deng et al., 2024; Zheng et al., 2024; Lai et al., 2024; Pan et al., 2024) have garnered particular attention due to their extensive application prospects and unique potential for fostering authentic autonomous intelligence within the digital ecosystem.
@@
Challenges. In this work, we propose to train high-performance web agents based on open LLMs within online environments, specifically utilizing WebArena (Zhou et al., 2023a). Our investigation has identified several critical challenges inherent to this task: 1) Insufficiency of training tasks: In contrast to offline datasets (Deng et al., 2024; Rawles et al., 2024) that facilitate agent training and evaluation on human-annotated oracle trajectories, online benchmarks such as WebArena typically provide only a limited test set for evaluation purposes. This dearth of predefined training tasks significantly impedes the effective training of agents within these environments. 2) Sparsity and cost of feedback signals: The assessment of success for arbitrary web browsing tasks is difficult in the absence of task-specific evaluation functions. Moreover, unlike tasks in certain GUI datasets (e.g., AITW (Rawles et al., 2024) and WebShop (Yao et al., 2022)), those in WebArena are typically of long horizons, with oracle solutions averaging about 10 steps. This characteristic introduces substantial sparsity in the available signals during online exploration. 3) Policy distribution drift in online learning: The absence of a predefined training set necessitates online exploration, inevitably leading to distribution drift in the agent’s policy. This phenomenon is likely to induce catastrophic forgetting and performance degradation over time.
The WebRL Framework. In response to these challenges, we introduce WebRL, a self-evolving online curriculum reinforcement learning framework designed for training LLM web agents. To the best of our knowledge, this represents the first systematic framework enabling effective reinforcement learning for LLM web agents from initialization in online web environments. Through the application of WebRL, we have successfully transformed a Llama-3.1-8B model into a proficient LLM web agent, elevating its success rate (SR) on WebArena-Lite (Zhou et al., 2023a; Liu et al., 2024) from an initial 4.8% to 42.4% across a diverse set of five websites. Furthermore, when applied to Llama-3.1-70B, we achieve a remarkable 49.1% SR, significantly surpassing the performance of the most advanced proprietary LLM API (GPT-4-Turbo, 17.6% SR) and the previous state-of-the-art web agents trained on open-source LLMs (AutoWebGLM (Lai et al., 2024), 18.2% SR).
@@ 2 WebRL: Self-Evolving Online Curriculum RL
We present a self-evolving online curriculum learning framework designed for training web agents, targeting the WebArena (Zhou et al., 2023a) environment. In this system, as illustrated in Figure 2, the agent continuously interacts with its environment to collect real-time trajectory data. This interaction is guided by the self-evolving curriculum learning strategy that dynamically generates tasks, effectively mitigating the insufficiency of training tasks. Furthermore, the tasks generated by the self-evolving curriculum learning strategy are tailored to the agent’s current proficiency, thereby increasing the likelihood of receiving positive feedback and alleviating the challenge of sparse feedback signals. Additionally, we train an outcome-supervised reward model (ORM) to evaluate task success. We introduce a KL-constrained policy update algorithm that prevents severe policy shifts during curriculum learning. A replay buffer is also utilized to retain prior knowledge and mitigate the risks of catastrophic forgetting. These techniques enable the agent to improve incrementally, progressively handling more complex tasks.
@@
ORM Training. In the curriculum learning process, we need to determine whether the corresponding instruction is completed based on the trajectory generated by the agent. Due to the lack of feedback from the environment, we train an LLM as the outcome-supervised reward model $\mathcal{M}_{\text{ORM}}$ to automate this task success evaluation. $\mathcal{M}_{\text{ORM}}$ allows us to assess the agent’s rollout trajectory for any given task, providing a binary reward signal (0 for failure and 1 for success).
@@ 2.1 Reinforcement Learning for LLMs in Online Web Environments
A typical challenge in training LLM web agents within WebArena is the scarcity of training tasks, resonating with the situation of developing real-world web agents. Although the recent work (Liu et al., 2024) has curated a trajectory fine-tuning set for WebArena, it only contains around 1k instructions with oracle trajectories, far from enough for training strong LLM web agents.
Therefore, we set our study to an online curriculum learning setting, where the model progressively encounters and learns a new set of tasks at each phase in the process to improve data efficiency. Considering the setting, a major challenge here is to avoid excessive policy distribution drift during each learning phase, which could lead to the catastrophic forgetting of previously acquired knowledge. Traditional approaches typically mitigate the issue by mixing data from different phases. However, in web agent tasks, intermediate steps do not receive direct process rewards, with only weak signals from the outcome of the final state. Consequently, even if an intermediate step is executed correctly, an error in later steps can easily lead to the final failure, resulting in misjudgment of the intermediate step and making it difficult to be reused. As a result, in this work, we primarily seek algorithmic improvements to address policy distribution drift more directly.
@@
$\beta\log\frac{\pi^{*}(a_{t}|s_{t},I)}{\pi_{\text{ref}}(a_{t}|s_{t},I)}=r(s_{t},a_{t},I)$ (4)
The work of (Ng et al., 1999) indicates that two rewards $r$ and $r^{\prime}$ are equal and will yield the same optimal policy if they satisfy $r^{\prime}(s_{t},a_{t})=r(s_{t},a_{t})+\phi(s_{t+1})-\phi(s_{t})$. Consequently, eq. 4 can be rewritten as:
@@
where $\nu(s)$ represents the distribution of experience in this phase.
@@
where $\lambda$ is a balancing factor that controls the trade-off between bias and variance in advantage estimation. We set $\lambda$ as 0.5 in our work.
Experience Replay Buffer with Actor Confidence Filtering. In addition to controlling the policy distribution drift at the algorithmic level through KL, we also implement an adaptive replay buffer to alleviate knowledge forgetting at the data level. Specifically, we only store those successful trajectories (which can be sparse) from each phase in the replay buffer. During phase $i$, we use the actor from the last phase to compute the perplexity of all actions in the buffer. Actions with a perplexity within the range of 1/0.95 to 1/0.5, along with their corresponding states, are added to the training data for the current phase. This filtering process excludes both over-familiar data and data that remains too challenging for the actor. Additionally, by storing only successful trajectories, we avoid the challenge of accurately estimating intermediate states for incorrect trajectories from previous phases.
2.2 Self-evolving New Instruction for Curriculum Learning
To achieve continuous improvement, we implement a self-evolving curriculum-learning strategy that generates instructions tailored to the agent’s current skill level. This strategy incrementally enhances the agent’s capabilities by progressively adapting and advancing the complexity of tasks in accordance with the agent’s development. In each phase, we implement a two-step process of generation and filtering, to produce tasks that are incrementally more challenging, while still being suitable for the agent’s current capability. During the generation phase, we use the in-breadth evolving approach (Xu et al., 2023) to create new instructions. We select instructions the model failed to complete in the previous interaction phase as seeds for generating new instructions. Detailed prompts are provided in the Appendix § C. To ensure that the generated instructions are both feasible in the target environment and aligned with the desired difficulty level, we first filter them using the trained critic. Specifically, we use the critic to evaluate each new instruction by considering its initial state. We select instructions with critic scores between 0.05 and 0.75, ensuring that only tasks meeting our difficulty criteria are retained. Then, we manually check the instructions to eliminate instructions that are clearly unreasonable or impossible to accomplish. The resulting set of instructions is used for interaction and training in the next phase.
@@ 3.1 Environments and Baselines
Environments. The effectiveness of WebRL and baseline methods is evaluated using the WebArena environment (Zhou et al., 2023a). WebArena is particularly well-suited to our needs, as it provides a highly interactive platform that supports online learning. Additionally, WebArena encompasses a variety of websites, including OpenStreetMap (Map), Reddit, GitLab, online store content management system (CMS), and OneStopShop (OSS), making it an ideal benchmark for comprehensively assessing model performance on web tasks. In the original WebArena environment, a total of 812 instructions are provided. Considering the cost of testing, we use 165 test cases from WebArena-Lite (Liu et al., 2024) for evaluation.
Baselines. We compare WebRL with proprietary LLMs utilizing prompting techniques, as well as open-sourced LLMs trained with alternative methods. For proprietary models, we select GPT-4-Turbo-2024-0409 (GPT-4-Turbo) (Achiam et al., 2023) and GPT-4o. In addition to AWM (Wang et al., 2024) and WebPilot (Zhang et al., 2024f), we also use the results of models under the simple prompt as baselines. Details of the simple prompt can be seen in Appendix § C. For the open-source models, in addition to using these models with the simple prompt as baselines, we also train Llama3.1 (Dubey et al., 2024) and GLM-4-9B (GLM et al., 2024) using various approaches as baselines. Specifically, we employ imitation learning, also referred to as supervised fine-tuning (SFT), to train these models. The training data is derived from publicly available human-labeled demonstrations, sourced from the WebArena-Lite. In addition, we also explore several reinforcement learning methods for comparison, including Filtered Behavior Cloning (Filtered BC) (Pan et al., 2024), advantage-weighted regression (AWR) (Peng et al., 2019) and DigiRL (Bai et al., 2024). For WebRL and the reinforcement learning-based baselines, we utilize the SFT-trained model as the initial model for the actor. The critic is similarly based on the SFT-trained model, with the addition of a randomly initialized value head. The training details of WebRL and baselines can be found in Appendix § A.
Table 1: Task success rate (SR) of WebRL and other comparison methods, evaluated on WebArena-Lite (Zhou et al., 2023a; Liu et al., 2024), a human-verified subset of WebArena (* denotes results on full WebArena taken from literature reporting). The best and second-best models are highlighted.
@@
ORM. WebArena-Lite (Liu et al., 2024) provides training samples along with a corresponding reward function. We further enhance this set of data by introducing task rewrites, as well as modifying certain data variables, such as place names and product names. We also make adjustments to the associated reward function. $\mathcal{M}_{\text{ORM}}$ is trained using rollouts of WebRL and part of baseline methods on this set of tasks, with evaluation results determined by the reward function. More details can be found in Appendix § A.
@@ 3.7 Ablation Study
We conduct an ablation study to evaluate the impact of the replay buffer, KL-constrained policy update algorithm, and the curriculum learning strategy on WebRL. To assess their contributions, we compare WebRL with four alternative models: (1) WebRL w/o replay buffer, where training uses only the current interaction trajectory, (2) WebRL w/o KL, where the policy is updated using SFT but retains the replay buffer, (3) WebRL w/o KL & replay buffer, which uses neither a replay buffer nor the KL-constrained policy update algorithm, and (4) DigiRL, which employ online learning on a fixed set of instructions and AWR algorithm to update the policy.
The results, shown in Figure 5, demonstrate that all the components used by WebRL are essential. (1) The role of the replay buffer. The results reveal that when the replay buffer is removed, both WebRL w/o replay buffer and WebRL w/o KL & replay buffer experience worsening performance over time. This decline occurs because the models lose access to earlier experiences and focus only on recent data, leading to knowledge degradation. (2) The role of the KL-constrained policy update algorithm. Comparing WebRL and WebRL w/o KL, WebRL consistently performs better, due to the incorporation of KL-constrained policy update algorithm. While both use the replay buffer, WebRL benefits from reinforcement learning, which allows it to extract valuable information even from failed trajectories, unlike imitation learning in SFT, which discards error trajectories entirely. When the replay buffer is not used, the KL-constrained policy update algorithm degrades more slowly than SFT because it better retains past knowledge by controlling KL divergence. In contrast, SFT quickly overfits the current phase’s data and consistently underperforms its initial value. Overall, the KL-constrained policy update algorithm is more effective at balancing the retention of past knowledge with the learning of new information. (3) The role of the self-evolving curriculum learning strategy. When comparing WebRL to DigiRL, both exhibit an overall upward trend due to online learning. However, DigiRL progresses more slowly and reaches a lower performance ceiling because it operates within a fixed task framework, whereas WebRL generates new tasks that adapt to its evolving capabilities. This highlights the effectiveness of our self-evolving curriculum learning approach.

### v2

@@ Abstract
Figure 1: (a) Compared with all proprietary and open-sourced LLMs, GLM-4-9B with WebRL achieves the best results. (b) The performance of GLM-4-9B on WebArena-Lite (Zhou et al., 2024a; Liu et al., 2024b), trained using WebRL, shows significant improvement over other baselines across all five evaluated websites.
@@ 1 Introduction
Large language models (LLMs) have exhibited not only superior comprehension of human language, commonsense reasoning, and knowledge acquisition, but also significant potential in complex planning and logical reasoning, indicating their promising trajectory towards serving as autonomous LLM agents (Wang et al., 2024a; Liu et al., 2024a). A diverse array of applications for LLM agents has proliferated, encompassing domains such as code generation (Jimenez et al., 2024), database manipulation (Zhou et al., 2023; Gu et al., 2024), and graphical user interface (GUI) interaction (Rawles et al., 2024; Yang et al., 2023; Xie et al., 2024). Among these, web agents powered by LLMs (Deng et al., 2024; Zheng et al., 2024; Lai et al., 2024; Pan et al., 2024) have garnered particular attention due to their extensive application prospects and unique potential for fostering authentic autonomous intelligence within the digital ecosystem.
@@
Challenges. In this work, we propose to train high-performance web agents based on open LLMs within online environments, specifically utilizing WebArena (Zhou et al., 2024a). Our investigation has identified several critical challenges inherent to this task: 1) Insufficiency of training tasks: In contrast to offline datasets (Deng et al., 2024; Rawles et al., 2024) that facilitate agent training and evaluation on human-annotated oracle trajectories, online benchmarks such as WebArena typically provide only a limited test set for evaluation purposes. This dearth of predefined training tasks significantly impedes the effective training of agents within these environments. 2) Sparsity and cost of feedback signals: The assessment of success for arbitrary web browsing tasks is difficult in the absence of task-specific evaluation functions. Moreover, unlike tasks in certain GUI datasets (e.g., AITW (Rawles et al., 2024) and WebShop (Yao et al., 2022)), those in WebArena are typically of long horizons, with oracle solutions averaging about 10 steps. This characteristic introduces substantial sparsity in the available signals during online exploration. 3) Policy distribution drift in online learning: The absence of a predefined training set necessitates online exploration, inevitably leading to distribution drift in the agent’s policy. This phenomenon is likely to induce catastrophic forgetting and performance degradation over time.
The WebRL Framework. In response to these challenges, we introduce WebRL, a self-evolving online curriculum reinforcement learning framework designed for training LLM web agents. To the best of our knowledge, this represents the first systematic framework enabling effective reinforcement learning for LLM web agents from initialization in online web environments. Through the application of WebRL, we have successfully transformed a Llama-3.1-8B model into a proficient LLM web agent, elevating its success rate (SR) on WebArena-Lite (Zhou et al., 2024a; Liu et al., 2024b) from an initial 4.8% to 42.4% across a diverse set of five websites. Furthermore, when applied to Llama-3.1-70B, we achieve a remarkable 49.1% SR, significantly surpassing the performance of the most advanced proprietary LLM API (GPT-4-Turbo, 17.6% SR) and the previous state-of-the-art web agents trained on open-source LLMs (AutoWebGLM (Lai et al., 2024), 18.2% SR).
@@ 2 WebRL: Self-Evolving Online Curriculum RL
We present a self-evolving online curriculum learning framework designed for training web agents, targeting the WebArena (Zhou et al., 2024a) environment. In this system, as illustrated in Figure 2, the agent continuously interacts with its environment to collect real-time trajectory data. This interaction is guided by the self-evolving curriculum learning strategy that dynamically generates tasks, effectively mitigating the insufficiency of training tasks. Furthermore, the tasks generated by the self-evolving curriculum learning strategy are tailored to the agent’s current proficiency, thereby increasing the likelihood of receiving positive feedback and alleviating the challenge of sparse feedback signals. Additionally, we train an outcome-supervised reward model (ORM) to evaluate task success. We introduce a KL-constrained policy update algorithm that prevents severe policy shifts during curriculum learning. A replay buffer is also utilized to retain prior knowledge and mitigate the risks of catastrophic forgetting. These techniques enable the agent to improve incrementally, progressively handling more complex tasks. The overall training process can be found in Algorithm 1.
@@
ORM Training. In the curriculum learning process, we need to determine whether the corresponding instruction is completed based on the trajectory generated by the agent. Due to the lack of feedback from the environment, we train an LLM as the outcome-supervised reward model $\mathcal{M}_{\text{ORM}}$ to achieve this task success evaluation. $\mathcal{M}_{\text{ORM}}$ is utilized to assess whether the agent’s rollout trajectory accomplishes a given task, providing a binary reward signal (0 for failure and 1 for success).
@@
2.1 Self-evolving New Instruction for Curriculum Learning
A typical challenge in training LLM web agents within WebArena is the scarcity of training tasks, resonating with the situation of developing real-world web agents. Although the recent work (Liu et al., 2024b) has curated a trajectory fine-tuning set for WebArena, it only contains around 1k instructions with oracle trajectories, far from enough for training strong LLM web agents. To address this limitation and drive continuous improvement, we employ a self-evolving curriculum learning strategy. This method generates new training instructions at each phase. As the phase progresses, the generated instructions become increasingly complex, allowing the agent’s capabilities to improve gradually. We implement a two-step process of generation and filtering, to produce tasks that are incrementally more challenging, while still being suitable for the agent’s current capability. During the generation step, we use the in-breadth evolving approach (Xu et al., 2023) to create new instructions. We select instructions the model failed to complete in previous interaction phases as seeds for generating new instructions. Detailed prompts are provided in the Appendix § D. To ensure that the generated instructions are both feasible in the target environment and aligned with the desired difficulty level, we first filter them using the trained critic. Specifically, we use the critic to evaluate each new instruction by considering its initial state. We select instructions with critic scores between 0.05 and 0.75, ensuring that only tasks meeting our difficulty criteria are retained. We manually review generated tasks and identify tasks that cannot be completed in WebArena. Based on these findings, we develop a prompt (Figure 22) and use GPT-4o to exclude infeasible tasks in WebArena automatically. The resulting set of instructions is used for interaction and training in this phase.
@@ 2.2 Reinforcement Learning for LLMs in Online Web Environments
In each phase of curriculum learning, the model progressively encounters and learns a new set of tasks. Considering this setting, a major challenge here is to avoid excessive policy distribution drift during each learning phase, which could lead to the catastrophic forgetting of previously acquired knowledge. Traditional approaches typically mitigate the issue by mixing data from different phases. However, in web agent tasks, intermediate steps do not receive direct process rewards, with only weak signals from the outcome of the final state. Consequently, even if an intermediate step is executed correctly, an error in later steps can easily lead to the final failure, resulting in misjudgment of the intermediate step and making it difficult to be reused. As a result, in this work, we primarily seek algorithmic improvements to address policy distribution drift more directly.
@@
$\beta\log\frac{\pi^{*}(a_{t}|s_{t},I)}{\pi_{\text{ref}}(a_{t}|s_{t},I)}=r(s_{t},a_{t},I)+V^{*}(s_{t+1},I)-V^{*}(s_{t},I)=A^{*}(s_{t},a_{t},I)$ (4)
@@
where $\nu(s)$ represents the distribution of experience in this phase. Note that our algorithm operates in the off-policy manner. A comprehensive derivation, thorough analysis, and a detailed comparison with other RL algorithms can be found in Appendix A.1.
@@
where $\lambda$ is a balancing factor that controls the trade-off between bias and variance in advantage estimation. We set $\lambda$ as 0.5 in our work. We approximate the true advantage function $A^{*}$ using the estimated advantage function $A$, which is derived from the value network $V$. The feasibility of this approximation is demonstrated in Appendix A.2, where we show that using $A$ to update the policy can lead to policy improvement.
Experience Replay Buffer with Actor Confidence Filtering. In addition to controlling the policy distribution drift at the algorithmic level through KL divergence, we also implement an adaptive replay buffer to alleviate knowledge forgetting at the data level. Specifically, we only store those successful trajectories (which can be sparse) from each phase in the replay buffer. During phase $i$, we use the actor from the last phase to compute the perplexity of all actions in the buffer. Actions with a perplexity within the range of 1/0.95 to 1/0.5, along with their corresponding states, are added to the training data for the current phase. This filtering process excludes both over-familiar data and data that remains too challenging for the actor. Additionally, by storing only successful trajectories, we avoid the challenge of accurately estimating intermediate states for incorrect trajectories from previous phases.
@@ 3.1 Environments and Baselines
Environments. The effectiveness of WebRL and baseline methods is evaluated using the WebArena environment (Zhou et al., 2024a). WebArena is particularly well-suited to our needs, as it provides a highly interactive platform that supports online learning. Additionally, WebArena encompasses a variety of websites, including OpenStreetMap (Map), Reddit, GitLab, online store content management system (CMS), and OneStopShop (OSS), making it an ideal benchmark for comprehensively assessing model performance on web tasks. In the original WebArena environment, a total of 812 instructions are provided. Considering the cost of testing, we use 165 test cases from WebArena-Lite (Liu et al., 2024b) for evaluation.
Baselines. We compare WebRL with proprietary LLMs utilizing prompting techniques, as well as open-sourced LLMs trained with alternative methods. For proprietary models, we select GPT-4-Turbo-2024-0409 (GPT-4-Turbo) (Achiam et al., 2023) and GPT-4o. In addition to AWM (Wang et al., 2024b) and WebPilot (Zhang et al., 2024f), we also use the results of models under the simple prompt as baselines. Details of the simple prompt can be seen in Appendix § D. For the open-source models, in addition to using these models with the simple prompt as baselines, we also train Llama3.1 (Dubey et al., 2024) and GLM-4-9B (GLM et al., 2024) using various approaches as baselines. Specifically, we employ imitation learning, also referred to as supervised fine-tuning (SFT), to train these models. The training data is derived from publicly available human-labeled demonstrations, sourced from the WebArena-Lite. In addition, we also explore several reinforcement learning methods for comparison, including Filtered Behavior Cloning (Filtered BC) (Pan et al., 2024), advantage-weighted regression (AWR) (Peng et al., 2019) and DigiRL (Bai et al., 2024). For WebRL and the reinforcement learning-based baselines, we utilize the SFT-trained model as the initial model for the actor. The critic is similarly based on the SFT-trained model, with the addition of a randomly initialized value head. The training details of WebRL and baselines can be found in Appendix § B.
Table 1: Task success rate (SR) of WebRL and other comparison methods, evaluated on WebArena-Lite (Zhou et al., 2024a; Liu et al., 2024b), a human-verified subset of WebArena (* denotes results on full WebArena taken from literature reporting). The best and second-best models are highlighted.
@@
ORM. WebArena-Lite (Liu et al., 2024b) provides training samples along with a corresponding reward function. We further enhance this set of data by introducing task rewrites, as well as modifying certain data variables, such as place names and product names. We also make adjustments to the associated reward function. $\mathcal{M}_{\text{ORM}}$ is trained using rollouts of WebRL and part of baseline methods on this set of tasks, with evaluation results determined by the reward function. More details can be found in Appendix § B.
@@ 3.7 Ablation Study
We conduct an ablation study to evaluate the impact of the replay buffer, KL-constrained policy update algorithm, and the curriculum learning strategy on WebRL. To assess their contributions, we compare WebRL with four alternative models: (1) WebRL w/o replay buffer, where training uses only the current interaction trajectory, (2) WebRL w/o KL, where the policy is updated using REINFORCE with value function baseline (the gradient is $\mathbb{E}_{\nu}[(A(s,a,I)\nabla_{\theta}\log\pi_{\theta}(a|s,I)]$) but retains the replay buffer, (3) WebRL w/o KL & replay buffer, which uses neither a replay buffer nor the KL-constrained policy update algorithm, and (4) WebRL w/o CL, which ablates the curriculum learning approach, utilizing only the instructions generated in the first phase.
The results, shown in Figure 5, demonstrate that all the components used by WebRL are essential. (1) The role of the replay buffer. The results reveal that when the replay buffer is removed, both WebRL w/o replay buffer and WebRL w/o KL & replay buffer experience worsening performance over time. This decline occurs because the models lose access to earlier experiences and focus only on recent data, leading to knowledge degradation. (2) The role of the KL-constrained policy update algorithm. Comparing WebRL and WebRL w/o KL, WebRL consistently performs better, due to the incorporation of KL-constrained policy update algorithm. When the replay buffer is not used, the KL-constrained policy update algorithm degrades more slowly than REINFORCE with a value function baseline because it better retains past knowledge by controlling KL divergence. In contrast, REINFORCE with a value function baseline quickly overfits the current phase’s data and consistently underperforms its initial value. Overall, the KL-constrained policy update algorithm is more effective at balancing the retention of past knowledge with the learning of new information. (3) The role of the self-evolving curriculum learning strategy. When comparing WebRL to WebRL w/o CL, both exhibit an overall upward trend due to online learning. However, WebRL w/o CL progresses more slowly and reaches a lower performance ceiling because it operates within a fixed task framework, whereas WebRL generates new tasks that adapt to its evolving capabilities. This highlights the effectiveness of our self-evolving curriculum learning approach. Additional ablation studies to further investigate the effects of curriculum learning are presented in Figure 14.